# Stable generative modeling using Schrödinger bridges

### v2

@@ 6. Numerical experiments
In this section, we illustrate our method through three numerical examples encompassing different ranges and focal points. In the first two examples, we generate samples using synthetic datasets with increasing dimensions. Our emphasis is on exploring the impact of different diffusions and of employing a variable bandwidth kernel. In the third example, we showcase the proposed conditional generative modeling in Section 5, applied to a stochastic subgrid-scale parametrization problem.
@@ 6.1. One-dimensional manifold
We begin by investigating the effect of the two noise models proposed, namely a constant diffusion as in (15) with constant bandwidth $K(x)=I$ and the case when the diffusion reproduces the sample covariance $C$ as in (18). We employ a Langevin sampler with the splitting scheme (20) with $\epsilon=0.009$ and (21), respectively. Figure 1 shows that choosing the sample covariance as the noise model is clearly advantageous. Whereas both noise models generate samples that reproduce the angular distribution the noise model using a constant diffusion is overdiffusive in the radial direction. In contrast the noise model using the sample covariance nicely reproduces the radial distribution.
We now investigate the effect of a variable bandwidth $K(x)$. We employ the noise model (21) with the sample covariance but use $K(x)$ to determine the diffusion map (cf. (9)). The Langevin sampler (21) is again initialized with the coordinates of the data point corresponding to the smallest angle in the data-sparse tail. Figure 2 (left) shows the samples projected onto the convex hull of the data, i.e. outputs of step (21b), when a uniform bandwidth $K(x)=I$ with $\epsilon=0.009$ is employed. Although the mean behaviour is well reproduced, it is seen that the generative model fails near the data-sparse tails for large and small angles. Here the value of $\epsilon$ is too small to generate significant diffusion and the samples are aligned on the (linear) convex hull of the widely separated data samples. To mitigate against this behaviour we employ a variable bandwidth $\rho(x)=\pi^{\beta}$ with $\beta=-1/5$ and a kernel density estimate $\pi(x)$. Figure 2 (right) shows how the variable bandwidth kernel is able to better reproduce the sampling in the data-sparse tail regions. Figure 3 shows the empirical histograms for the radius and the angle variables of the noisy samples corresponding to Figure 2 (i.e. outputs of step (21a)). Whereas both, the uniform and the variable bandwidth kernels, reproduce the radial distribution very well, the uniform bandwidth fails to reproduce the angular distribution in the tails where the diffusion is not sufficiently strong to let the sampler escape the convex hull of the data.
@@
The numerical experiments above suggest that we employ a noise model using the sample covariance $C$ combined with a variable bandwidth $K(x)$ to control eventual data sparse regions. When employing a variable bandwidth our method contains two hyperparameters which require tuning: the bandwidth factor $\epsilon$ and the exponent $\beta$ in the arbitrary choice of the variable bandwidth $\rho(x)$. Their role will be explored in the following subsection.
@@ 6.2. Multi-dimensional manifolds
subject to the constraint that
$P1_{M_{\mathrm{r}}}$ $=\frac{1}{M_{\mathrm{g}}}\left(1,\cdots,1\right)^{\top}\in\mathbb{R}^{M_{g}}$
$P^{\top}1_{M_{\mathrm{g}}}$ $=\frac{1}{M_{\mathrm{r}}}\left(1,\cdots,1\right)^{\top}\in\mathbb{R}^{M_{r}},$
@@ 6.3. Stochastic subgrid-scale parametrization
where $\varepsilon<1$ denotes the time scale separation between the slow resolved variables of interest $z\in{\mathbb{R}}^{d_{s}}$ and the unresolved fast degrees $y\in{\mathbb{R}}^{d_{f}}$. Note the notational difference between the time scale separation parameter $\varepsilon$ and the bandwidth parameter $\epsilon$ used to define the diffusion map. For $\varepsilon\ll 1$ the system is stiff and to ensure numerical stability a small time step $\Delta t<\varepsilon$ is needed. This, together with the potential high-dimensionality of the fast subspace $d_{f}\gg d_{s}$, constitutes a computational barrier for simulating the dynamics on the slow time scale of interest. Hence one is interested in obtaining an effective evolution equation for the slow resolved variables $z$ only which captures the essential effect of the unresolved variables $y$. Hence, we seek to determine the effective reduced dynamics
@@
Here $F_{z}(z)$ denotes a deterministic drift which we assume to be known a priori, possibly based on physical reasoning. The term $\psi(z)$ denotes the unknown closure term to be learned which may be deterministic or stochastic, and which parametrizes the unknown unresolved fast processes. Deterministic machine learning methods have previously been used to learn $\psi(z)$ as the average effect of the unresolved variables, i.e. the average of $g(z,y;\varepsilon)$ over the (conditional) invariant measure of the fast process [19, 20]. Deterministic maps, however, are not able to capture the resolved dynamics with sufficient statistical accuracy, and it is by now well established that the effective equation is often of a stochastic nature [37, 18, 25, 17]. In the case of infinite time scale separation there are explicit expressions for the effective drift and diffusion term of the effective slow dynamics. However, these terms include integrals over auto-correlation functions and are notoriously hard to estimate. Instead we propose to learn the closure term $\psi(z)$ and generate realisations $\psi(z)$ on the fly employing the conditional sampling algorithm described in Section 5. We consider the situation in which scientists have a good understanding of the resolved dynamics and know the slow vector field $F_{z}(z)$. Given data of the resolved variables $\{z_{n}\}_{n=1}^{N}$ sampled at equidistant times $\Delta t$, the closure term $\psi(z)$ capturing the effect of the unresolved dynamics (29b) can then be estimated as
$\psi(z_{n})=z_{n+1}-z_{n}-F_{z}(z_{n})\,\Delta t,$
for $n=1,\cdots,M-1$. Note that the closure term $\psi(z)$ typically includes effective diffusion as well as a correction to the drift term $F_{z}(z)$ [16]. The effective dynamics is then provided by the discrete stochastic surrogate model
(30) $z_{n+1}=z_{n}+F_{z}(z_{n})\,\Delta t+\psi_{n},$
where the subgrid-scale terms $\psi_{n}=\psi(z_{n})$ are generated as follows: Given a kernel $t_{i}(x)$ with $x=(z,\psi(z))^{T}$ and a $v_{\epsilon}$ obtained by the Sinkhorn algorithm as described in Section 3, and the $2d_{s}\times M$ data matrix $\mathcal{X}$ consisting of samples $\{x^{(j)}\}_{j=1}^{M}$, we perform for $j=1,\cdots,n_{s}$ with $n_{s}=100$ a discrete Langevin sampler for $x$
(31a) $z^{\ast}$ $=z_{n}$
(31b) $x_{1,j}$ $=z^{\ast}$
(31c) $x_{j+1/2}$ $=x_{j}+\sqrt{2\epsilon}\,\Xi_{j}$
(31d) $x_{j+1}$ $=\mathcal{Z}p_{\epsilon}(x_{j+1/2}),$
with $\Xi_{n}\sim{\rm N}(0,I)$. The first assignment $x_{1,n}=z^{\ast}=z_{n}$ ensures the conditioning of $\psi(z_{n})$ on $z^{\ast}=z_{n}$, and $n_{s}=100$ ensures that the generated samples $\psi(z_{n})$ are close to independent. We choose a fixed bandwidth with $\epsilon=0.001$.
@@
We look at the case of effective additive noise with $h(z)=1$ which does not require conditioning on the slow variable $z$, as well as at the case of multiplicative noise with $h(z)=z$. We used MATLAB’s built-in ode45 routine [35] to generate the time series with a time-scale separation parameter of $\varepsilon=0.01$. The time series is subsequently sub-sampled with $\Delta t=0.1$. Figure 8 shows a comparison between the full multi-scale system (32) - (33) with additive noise $h(z)=1$ and the stochastic surrogate model (30). The slow $z$-dynamics exhibits stochastic bimodal dynamics. Figure 9 shows a comparison for the multiplicative case $h(z)=z$ which yields unimodal slow dynamics. It is clearly seen that the surrogate model (30) obtained by the generative conditional sampler generates statistically reliable dynamics.
@@
Figure 8. Results for the stochastic subgrid-scale parametrization for the multi-scale system (32) - (33) with $\varepsilon=0.01$ and with additive noise $h(z)=1$. Shown are results obtained by integrating the full multi-scale system and by the stochastic subgrid-scale parametrization scheme using our generative sampler (31) trained with $M=120,000$. Left: Empirical histograms of the closure term $\psi=\psi(z)$. Middle: Time series of the slow variable $z(t)$. Right: Empirical histograms of the slow variable $z$.
@@
Figure 9. Results for the stochastic subgrid-scale parametrization for the multi-scale system (32) - (33) with $\varepsilon=0.01$ and with multiplicative noise $h(z)=z$. Shown are results obtained by integrating the full multi-scale system and by the stochastic subgrid-scale parametrization scheme using our generative sampler (31), trained with $M=20,000$. Left: Empirical histograms of the closure term $\psi=\psi(z)$. Middle: Time series of the slow variable $z(t)$. Right: Empirical histograms of the slow variable $z$.
@@ 7. Summary and outlook
We introduced a Schrödinger bridge based Langevin scheme for generative modeling. Our method combines sample-based Markov chain approximations with discrete-time Langevin-type sampling, yielding a nonparametric generative model that is unconditionally stable and geometrically ergodic. We showed numerically that employing a variable bandwidth kernel, in contrast to a fixed bandwidth kernel, results in generated samples with enhanced accuracy. The performance of the conditional generative model was showcased through its application to a stochastic subgrid-scale parametrization problem.
We considered here a uniform measure on the samples. If there is a change of measure due to, for example, observed data with a given likelihood, then the measure and the couplings can be appropriately modified to take into account the non-uniform measure.
Future research will delve into the theoretical foundations of the proposed scheme, including its convergence rate and scalability. Finally, the proposed methodology can be extended to score generative modeling by replacing all required score functions by Schrödinger bridge-based approximations and to interacting particle approximations of the Fokker–Planck equation, where the score function represents now diffusion [33].
@@ Acknowledgements.
This work has been funded by Deutsche Forschungsgemeinschaft (DFG) - Project-ID 318763901 - SFB1294. GAG acknowledges funding from the Australian Research Council, grant DP220100931. FL and YMM acknowledge support from the US Department of Energy, Office of Advanced Scientific Computing Research, award DE-SC0023188. We thank Ricardo Baptista for pointing out the relationship to the barycentric projection and Greg Pavliotis for pointing out the connection to the Klimontovitch stochastic integral formulation.

### v3

@@ 5. Numerical experiments
In this section, we illustrate our method through three numerical examples encompassing different ranges and focal points. In the first two examples, we generate samples using synthetic datasets with increasing dimensions. Our emphasis is on exploring the impact of different diffusions and of employing a variable bandwidth kernel. In the third example, we showcase the proposed conditional generative modeling in Section 4, applied to a stochastic subgrid-scale parametrization problem. In the final example in Section 5.4, we employ conditional sampling to generate trajectories of the Lorenz-63 system [33] from previously computed training samples.
@@ 5.1. One-dimensional manifold
We begin by investigating the effect of the two noise models proposed, namely a constant diffusion as in (20) with constant bandwidth $K(x)=I$ and the case when the diffusion reproduces the sample covariance $C$ as in (23). In both cases we use a constant bandwidth $K=I$ in (9) and (14) when forming the Schrödinger bridge. We employ a Langevin sampler with the splitting scheme (29) with $\epsilon=0.009$ and (30), respectively. Figure 2 shows that choosing the sample covariance as the noise model is clearly advantageous. Whereas both noise models generate samples that reproduce the angular distribution the noise model using a constant diffusion is overdiffusive in the radial direction. In contrast the noise model using the sample covariance nicely reproduces the radial distribution.
We now investigate the effect of a variable bandwidth $K(x)$. We employ the noise model (30) with the sample covariance but use $K(x)$ to determine the diffusion map (cf. (14)). The Langevin sampler (30) is again initialized with the coordinates of the data point corresponding to the smallest angle in the data-sparse tail. Figure 3 (left) shows the samples projected onto the convex hull of the data, i.e. outputs of step (30b), when a uniform bandwidth $K(x)=I$ with $\epsilon=0.009$ is employed. Although the mean behaviour is well reproduced, it is seen that the generative model fails near the data-sparse tails for large and small angles. Here the value of $\epsilon$ is too small to generate significant diffusion and the samples are aligned on the (linear) convex hull of the widely separated data samples. To mitigate against this behaviour we employ a variable bandwidth $\rho(x)=\pi^{\beta}$ with $\beta=-1/5$ and a kernel density estimate $\pi(x)$. Figure 3 (right) shows how the variable bandwidth kernel better reproduces the distribution in the data-sparse tail regions. Figure 4 shows the empirical histograms for the radius and the angle variables of the noisy samples corresponding to Figure 3 (i.e. outputs of step (30a)). Whereas both, the uniform and the variable bandwidth kernels, reproduce the radial distribution very well, the uniform bandwidth fails to reproduce the angular distribution in the tails where the diffusion is not sufficiently strong to allow for efficient mixing and the diffusion process gets stuck in the data sparse region.
@@
The numerical experiments above suggest that we employ a noise model using the sample covariance $C$ combined with a variable bandwidth $K(x)$ to control eventual data sparse regions. When employing a variable bandwidth our method contains two hyper-parameters which require tuning: the bandwidth factor $\epsilon$ and the exponent $\beta$ in the arbitrary choice of the variable bandwidth $\rho(x)$. Their role will be explored in the following subsection.
@@ 5.2. Multi-dimensional manifolds
subject to the constraints that
$\sum_{j=1}^{M_{\mathrm{r}}}P_{ij}=M_{\mathrm{g}}^{-1},\qquad\sum_{i=1}^{M_{\mathrm{g}}}P_{ij}=M_{\mathrm{r}}^{-1},$
@@ 5.3. Stochastic subgrid-scale parametrization
where $\varepsilon<1$ denotes the time scale separation between the slow resolved variables of interest $z\in{\mathbb{R}}^{d_{s}}$ and the unresolved fast degrees $y\in{\mathbb{R}}^{d_{f}}$. Note the notational difference between the time scale separation parameter $\varepsilon$ and the bandwidth parameter $\epsilon$ used to define the diffusion map. For $\varepsilon\ll 1$ the system is stiff and to ensure numerical stability a small time step $\Delta t<\varepsilon$ is needed. This, together with the potential high-dimensionality of the fast subspace $d_{f}\gg d_{s}$, constitutes a computational barrier for simulating the dynamics (50a)–(50b) on the slow time scale of interest. Hence one is interested in obtaining an effective evolution equation for the slow resolved variables $z$ only which captures the essential effect of the unresolved variables $y$. Hence, we seek to determine the effective reduced dynamics
@@
Here $F_{z}(z)$ denotes a deterministic drift which we assume to be known a priori, possibly based on physical reasoning. The term $\psi(z)$ denotes the unknown closure term to be learned which may be deterministic or stochastic, and which parametrizes the unknown unresolved fast processes. Deterministic machine learning methods have previously been used to learn $\psi(z)$ as the average effect of the unresolved variables, i.e. the average of $g(z,y;\varepsilon)$ over the (conditional) invariant measure of the fast process [20, 21]. Deterministic maps, however, are not able to capture the resolved dynamics with sufficient statistical accuracy, and it is by now well established that the effective equation is often of a stochastic nature [38, 19, 27, 18]. In the case of infinite time scale separation there are explicit expressions for the effective drift and diffusion term of the effective slow dynamics. However, these terms include integrals over auto-correlation functions and are notoriously hard to estimate. Instead we propose to learn the closure term $\psi(z)$ and generate realisations $\psi(z)$ on the fly employing the conditional sampling algorithm described in Section 4. We consider the situation in which scientists have a good understanding of the resolved dynamics and know the slow vector field $F_{z}(z)$. Given data of the resolved variables $\{z^{(i)}\}_{i=0}^{M}$ sampled at equidistant times $\Delta t$, the associated closure term $\psi^{(i)}$ capturing the effect of the unresolved dynamics (50b) can then be estimated as
$\psi^{(i)}:=z^{(i)}-z^{(i-1)}-F_{z}(z^{(i-1)})\,\Delta t,$
for $i=1,\cdots,M$. This produces the training samples $\{x^{(i)}=(z^{(i-1)},\psi^{(i)})\}_{i=1}^{M}$ which we collect in the $2d_{s}\times M$ data matrix $\mathcal{X}$.
The effective dynamics is then provided by the discrete stochastic surrogate model
(51) $z_{k}=z_{k-1}+F_{z}(z_{k-1})\,\Delta t+\psi_{k}$
for $k\geq 1$ and given $z_{0}$, where the subgrid-scale terms $\psi_{k}$ are generated as follows: Given the conditional mean $m(x;\epsilon)$ as described in Section 3, we perform the discrete Langevin sampler
(52a) $\hat{X}_{n}$ $=(z_{k-1},P_{2}X_{n})$
(52b) $X_{n+1/2}$ $=\hat{X}_{n}+\sqrt{2}\,\Xi_{n}$
(52c) $X_{n+1}$ $=m(X_{n+1/2};\epsilon),$
for $n=1,\cdots,n_{s}$ with $n_{s}=100$ and $\Xi_{n}\sim{\rm N}(0,\epsilon I)$. Here $P_{2}:\mathbb{R}^{2d_{s}}\to\mathbb{R}^{d_{s}}$ is a projector with $P_{2}x=\psi$. We finally set $\psi_{k}=P_{2}X_{n_{s}}$. The assignment of the first component of $\hat{X}_{n}$ to $z_{k-1}$ ensures the conditioning of $\psi_{k}$ on $z_{k-1}$, and $n_{s}=100$ ensures that the generated samples $\psi_{k}$ are close to independent from $\psi_{k-1}$. We choose a fixed bandwidth with $\epsilon=0.001$.
@@
We look at the case of effective additive noise with $h(z)=1$ which does not require conditioning on the slow variable $z$, as well as at the case of multiplicative noise with $h(z)=z$. We used MATLAB’s built-in ode45 routine [36] to generate the time series with a time-scale separation parameter of $\varepsilon=0.01$. The time series is subsequently sub-sampled with $\Delta t=0.1$. Figure 9 shows a comparison between the full multi-scale system (53)–(54) with additive noise $h(z)=1$ and the stochastic surrogate model (51). The slow $z$-dynamics exhibits stochastic bimodal dynamics. Figure 10 shows a comparison for the multiplicative case $h(z)=z$ which yields unimodal slow dynamics. It is clearly seen that the surrogate model (51) obtained by the generative conditional sampler generates statistically reliable dynamics.
We remark that the above proposed method to draw random samples from the closure term does not require the knowledge of the time-scale separation parameter $\varepsilon$. The closure term $\psi$ only depends on the time series $\{z^{(i)}\}_{i=0}^{M}$ of the resolved variables and the knowledge of the slow vector field $F_{z}$ (which may be zero).
In the next subsection we show that one can indeed draw full sample trajectories of a dynamical system even without any time-scale separation using the conditional Schrödinger bridge sampler.
@@
Figure 9. Results for the stochastic subgrid-scale parametrization for the multi-scale system (53)–(54) with $\varepsilon=0.01$ and with additive noise $h(z)=1$. Shown are results obtained by integrating the full multi-scale system and by the stochastic subgrid-scale parametrization scheme using our generative sampler (52) trained with $M=120,000$. Left: Empirical histograms of the closure term $\psi=\psi(z)$. Middle: Time series of the slow variable $z(t)$. Right: Empirical histograms of the slow variable $z$.
@@
Figure 10. Results for the stochastic subgrid-scale parametrization for the multi-scale system (53)–(54) with $\varepsilon=0.01$ and with multiplicative noise $h(z)=z$. Shown are results obtained by integrating the full multi-scale system and by the stochastic subgrid-scale parametrization scheme using our generative sampler (52), trained with $M=20,000$. Left: Empirical histograms of the closure term $\psi=\psi(z)$. Middle: Time series of the slow variable $z(t)$. Right: Empirical histograms of the slow variable $z$.
5.4. Generative modelling of dynamical systems
We now employ the Schrödinger bridge sampler to perform sequential conditional sampling and generate typical trajectories from a dynamical system. We consider the Lorenz-63 system [33] for $y=(y_{1},y_{2},y_{3})$ given by (54) with $\varepsilon=1$. Given a long time trajectory $\{y^{(i)}\}_{i=0}^{M}$ with equidistantly sampled data points $y^{(i)}=y(t_{i})$ with $t_{i+1}-t_{i}=\Delta t$ for all $i$, we construct the Schrödinger bridge as a coupling for the joint probability function $\pi(y(t),y(t+\Delta t))$ from the $M$ training pairs $x^{(i)}=(y^{(i-1)},y^{(i)})\in\mathbb{R}^{6}$. Our aim is to generate new trajectories $\{y_{k}\}_{k\geq 0}$ in time intervals of $\Delta t$ for given $y_{0}=y(t_{0})$ and $y_{1}=y(t_{1})$ sequentially, similarly to the generative modelling of texts: Given $x_{k-1}=(y_{k-1},y_{k})$, generate a new $x_{k}=(y_{k},y_{k+1})$, which is not part of the initial training samples. Obviously the first component of $x_{k}$ needs to be conditioned on the second component of $x_{k-1}$. As in Section 5.3, to decorrelate we perform $n_{s}=20$ Langevin sampling steps
(55a) $\hat{X}_{n}$ $=(y_{k},P_{2}X_{n})$
(55b) $X_{n+1/2}$ $=\hat{X}_{n}+\sqrt{2}\,\Xi_{n}$
(55c) $X_{n+1}$ $=m(X_{n+1/2};\epsilon),$
for given $y_{k}$ with $\Xi_{n}\sim{\rm N}(0,\epsilon I)$. Here $P_{2}:\mathbb{R}^{6}\to\mathbb{R}^{3}$ denotes the project onto the second three components of $x\in\mathbb{R}^{6}$. We finally set $y_{k+1}=P_{2}X_{n_{s}}$.
In Figure 11 we show that the conditional sampling produces trajectories which resemble those of the actual Lorenz-63 system, as well as having the same asymptotic statistical behaviour as seen by the reconstruction of the famous butterfly attractor. We choose a fixed bandwidth with $\epsilon=0.05$, and used $10,000$ training data $x(t_{n})$ with $\Delta t=0.1$.
The proposed approach can be compared to a direct approximation of the time-$\Delta t$-propagator
(56) $y_{n+1}=\Psi(y_{n})$
via, for example, a random feature map approximation as proposed in [20]. More specifically, using a smaller observation interval of $\Delta t=0.02$ and noisy data, highly accurate predictions were achieved in [20] using $D_{r}=300$ randomly chosen feature maps. The output weights $W\in\mathbb{R}^{3\times D_{r}}$ were learned using the ensemble Kalman filter [45]. In comparison, the proposed Schrödinger bridge sampler requires a larger training set ($M=10,000$ versus $M=4,000$ in [20]), but involves only a single tuning parameter $\epsilon$ and repeated computation of (17).
Figure 11. Comparison of generated data from the Schrödinger bridge sampler with the original dynamics of the Lorenz-63 system (54). Left: Typical trajectories of $y_{1}$ obtained from a simulation of (54) (top) and the generated data using conditional sampling (55) (bottom). Right: The corresponding attractor for the simulated and generated data.
@@ 6. Summary and outlook
We have introduced a Schrödinger bridge based Langevin scheme for generative modeling. Our method combines sample-based Markov chain approximations with discrete-time Langevin-type sampling, yielding a non-parametric generative model that is unconditionally stable and geometrically ergodic. Even though the approach is entirely data-driven, the generative model possesses a single tuning parameter, which is the choice of the step-size $\epsilon>0$. Practical choices of $\epsilon$ depend both on the number of training samples, $M$, and the properties of the target distribution $\pi(x)$ on $\mathbb{R}^{d}$. We showed numerically that employing a variable bandwidth kernel, in contrast to a fixed bandwidth kernel, results in generated samples with enhanced accuracy. However, employing variable a bandthwith kernel comes with an added layer of complexity and requires delicate tuning.
In terms of practical applications, the performance of the conditional generative model was showcased through its applications to a stochastic subgrid-scale parametrization problem and the generation of trajectories of the chaotic Lorenz-63 system. In both cases, the conditional sampler relied on the availability of a single simulated trajectory, which served as the training samples. A constant bandwith kernel has been employed in all these examples.
We considered in this paper only equally weighted training samples $x^{(i)}\sim\pi(x)$. If there is a change of measure due to, for example, observed data with a given likelihood, then the measure and the couplings can be appropriately modified to take into account the non-uniform measure, i.e, weighted training samples.
Furthermore, all computational examples considered in this paper were low-dimensional, which is intrinsic to Schrödinger bridge approximations of the semi-group $\exp(\epsilon\mathcal{L})$ [58]. This could be seen as a major disadvantage in comparison to the sophisticated deep neural network architectures typically used in high-dimensional SGMs [23, 52]. A first step to close this gap has been taken in [22], where a localized Schrödinger bridge sampler has been proposed and implemented for high-dimensional generative sampling problems. Interestingly, and worth further investigation, the localized Schrödinger bridge sampler can be viewed as a variant of multi-head self attention [56, 48].
Additional future research will delve into the theoretical foundations of the proposed scheme, including its convergence rate and scalability extending results from [58]. Finally, the proposed methodology can be extended to score generative modeling by replacing all required score functions by Schrödinger bridge-based approximations and to interacting particle approximations of the Fokker–Planck equation, where the score function represents now diffusion [34].
@@ Acknowledgements.
This work has been funded by Deutsche Forschungsgemeinschaft (DFG) - Project-ID 318763901 - SFB1294. GAG acknowledges funding from the Australian Research Council, grant DP220100931. FL and YMM acknowledge support from the US Department of Energy, Office of Advanced Scientific Computing Research, award DE-SC0023188. We thank Ricardo Baptista for pointing out the relationship to the barycentric projection.